OpenStudy (anonymous):

How to graph -2x = 2y + 5

OpenStudy (noorik.):

first you should make the function like that: 2y = -2x-5 then you start giving x positive and negative values to get y values, then draw these (x,y) points on the coordination graph

OpenStudy (shamil98):

Change to slope-intercept form, and then graph the line (plot the y-intercept and use the slope(the rise/run) to form the line).

OpenStudy (a1234):

Put it in slope-intercept form first. -2x = 2y + 5 -2y = 2x + 5 y = -x - 5/2 Plot the point -5/2. You see that the slope is -1. Go one down and one right from that point and keep doing so until you see the curve. Draw a line through them.

OpenStudy (anonymous):

ty

OpenStudy (a1234):

YW :)